# Strong Normalization for the Safe Fragment of a Minimal Rewrite System: A Triple-Lexicographic Proof and a Conjecture on the Unprovability of Full Termination for Any Relational Operator-Only TRS

### v3

@@ A.1 Lean kernel location
A.2 Triple-lexicographic measure and duplication handling (DM/MPO)
The measure $\mu^{3}(t)=(\delta\text{-flag}(t),\,\kappa^{M}(t),\,\mu_{\mathrm{ord}}(t))$ is ordered by lex:
@@ A.2 Triple-lexicographic measure and duplication handling (DM/MPO)
  • Ordinal $\mu_{\mathrm{ord}}$: resolves non-duplicating ties; right-addition is not strictly monotone; absorption $\alpha+\beta=\beta$ needs $\omega\leq\beta$.
@@
Aggregation for $\kappa^{M}$ (union, not sum)
It is emphasized that $\kappa^{M}$ aggregates via multiset union ($\cup$), not numeric addition. For duplicating rules, the multiset of piece-weights on the RHS is DM-smaller than the singleton multiset containing the LHS redex weight, yielding a strict drop in the $\kappa^{M}$ component. For non-duplicating rules, $\kappa^{M}$ ties by definitional equality and the ordinal $\mu_{\mathrm{ord}}$ resolves the branch. In particular, for unguarded instances of eqW-refl and merge-cancel a $\kappa$-branch strategy is employed: when $\kappa^{M}(a)\neq 0$ a left-lex drop is obtained via DM; when $\kappa^{M}(a)=0$ the $\kappa^{M}$ component ties by rfl and the strict decrease is witnessed in the $\mu_{\mathrm{ord}}$ coordinate (right branch of Prod.Lex).
@@
Figure 1: Triple-lex measure components. Duplicators decrease via $\kappa^{M}$; non-duplicating ties via $\mu_{\mathrm{ord}}$.
@@ A.3 Newman scope (guarded safe relation)
SN + local confluence implies confluence; in the artifact this is instantiated for the safe relation via an Acc-based star-star join. Scope note: all confluence statements and their Newman instantiations are for SafeStep only. Full Step is not locally joinable at root for eqW a a with $\kappa^{M}(a)=0$; accordingly full-step confluence is not claimed.
@@ A.4 Module map
[…]
Table 3: Per-rule decrease witnesses used in the SafeStep termination proof.
Rule Base component Precedence/Status Witness Source
merge void left $\mu$ (ordinal) (n/a) Theorem 4.1 Termination
merge void right $\mu$ (ordinal) (n/a) Theorem 4.2 Termination
merge cancel DM on $\kappa^{M}$ (redex >> pieces) head precedence Theorem 4.3 Termination
rec zero DM on $\kappa^{M}$ rec >> pieces Theorem 4.4 Termination
rec succ [^2] This row records the per-step decrease used in the SafeStep proof. It does not assert termination of the unrestricted full relation: rec_succ duplicates the step argument $s$ and is the structural source of the full-system termination barrier (cf. the conjecture/evidence discussion). $\delta$ phase bit (1 $\to$ 0) (n/a) Theorem 4.5 Termination
eqW refl MPO ($\mu$-first triple) head precedence Theorem 4.6 Termination
eqW diff MPO ($\mu$-first triple) head precedence Theorem 4.7 Termination
integrate (delta t) $\mu$ (unique target) (n/a) Theorem 4.8 Termination
Parenthetical aliases: for substitution convenience simp forms are exposed alongside the main rec lemmas.
@@ A.6 Local-join report (Confluence)
Root eqW a a, guard ne localJoin_eqW_refl_guard_ne refl blocked; only diff
Ctx merge a b (n-V, n-E) localJoin_ctx_merge_no_void_neq from root ‘merge’ lemma
@@
Ctx eqW a a, guard localJoin_ctx_eqW_refl_guard_ne refl blocked
Ctx eqW a a (norm) localJoin_ctx_eqW_refl_if_norm via normalizeSafe
@@ Appendix B Orientation Strategy
  • Non-duplicating rules use a $\mu$-first lex measure: triples $(\mu,\mathrm{sizeMPO},\delta)$ decrease directly via lemmas.
  • The rule rec_zero uses a DM multiset drop on $\kappa^{M}$.
  • The rule rec_succ drops the $\delta$ phase bit (1 $\to$ 0).
  • Merge-cancel decreases via DM on $\kappa^{M}$ under a base precedence where the redex strictly dominates its pieces.

### v4

@@ A.1 Lean kernel location
A.2 Auxiliary orientation sketch (DM/MPO)
This subsection records historical orientation notes used for boundary analysis. It is not the canonical SafeStep termination certificate in this paper, which is the computable stack in Meta/ComputableMeasure.lean. An auxiliary triple $\widetilde{\mu}^{3}(t)=(\delta\text{-flag}(t),\,\kappa^{M}(t),\,r_{\mathrm{aux}}(t))$ is ordered lexicographically:
@@ A.2 Auxiliary orientation sketch (DM/MPO)
  • Tie-break rank $r_{\mathrm{aux}}$: resolves non-duplicating ties when the first two coordinates are equal.
@@
Aggregation for $\kappa^{M}$
It is emphasized that $\kappa^{M}$ aggregates via multiset operations (union $\cup$ and addition ++), not scalar numeric addition. For duplicating rules, the multiset of piece-weights on the RHS is DM-smaller than the singleton multiset containing the LHS redex weight, yielding a strict drop in the $\kappa^{M}$ component. For non-duplicating rules, $\kappa^{M}$ ties by definitional equality and $r_{\mathrm{aux}}$ resolves the branch. In particular, for unguarded instances of eqW-refl and merge-cancel a $\kappa$-branch strategy is employed: when $\kappa^{M}(a)\neq 0$ a left-lex drop is obtained via DM; when $\kappa^{M}(a)=0$ the $\kappa^{M}$ component ties by rfl and the strict decrease is witnessed in the $r_{\mathrm{aux}}$ coordinate (right branch of Prod.Lex).
@@
Figure 1: Auxiliary triple-lex components. Duplicators decrease via $\kappa^{M}$; non-duplicating ties via $r_{\mathrm{aux}}$.
@@ A.3 Newman scope (guarded safe relation)
SN + local confluence implies confluence; in the artifact this is instantiated for the safe relation via an Acc-based star-star join. Scope note: all confluence statements and their Newman instantiations are for SafeStep only. Full Step has an explicit non-local-join witness at root, eqW void void; accordingly full-step confluence is not claimed.
@@ A.4 Module map
A.5 Auxiliary per-rule orientation (historical)
Earlier drafts used auxiliary orientation sketches for boundary exploration. They are historical context only and are not the canonical certified path of this paper.
The canonical per-rule decrease proofs used here are the computable lemmas in Meta/ComputableMeasure.lean (the drop_R_*_c family and measure_decreases_safe_c), with core SafeStep definitions in Meta/SafeStep_Core.lean.
@@ A.6 Local-join report (Confluence)
Root eqW a a, guard ne localJoin_eqW_refl_guard_ne both refl and diff blocked (vacuous join)
Ctx merge a b (n-V, n-E) localJoin_ctx_merge_no_void_neq from root merge lemma
@@
Ctx eqW a a, guard localJoin_ctx_eqW_refl_guard_ne both refl and diff blocked (vacuous join)
[…]
@@ Appendix B Orientation Strategy
  • All guarded rules are oriented in one computable lex order Lex3c over $(\delta\text{-flag},\kappa^{M},\tau)$.
  • The rule rec_succ is discharged by the outer $\delta\text{-flag}$ drop (1 $\to$ 0).
  • Multiset-sensitive branches (notably rec_zero) are discharged by DM drops on $\kappa^{M}$, while merge_cancel is discharged by $\tau$.
  • Tie branches are discharged by strict inequalities in the computable natural rank $\tau$.
  • Strong normalization follows from well-foundedness of Lex3c via inverse-image and yields wf_SafeStepRev_c.